OpenStudy (acespeedfighter):

1/4a = 2/3 . which of the following equals a in this equation? (32 points) O 1/6 O 3/8 O 11/12 O2 2/3

OpenStudy (calculusxy):

We need to use the division property of equality to solve this problem. So we need to divide both sides by 1/4 to isolate a and solve for it. \[\large \bf \frac{1}{4}a \div \frac{1}{4} = \frac{2}{3} \div \frac{1}{4}\] \[\large \bf \frac{2}{3} \div \frac{1}{4} = \frac{2}{3} \times \frac{4}{1} = \color{green}{\frac{8}{3}}\] Check: \[\large \bf \frac{1}{4} \times \frac{8}{3} = \frac{8}{12} = \color{red}{\frac{2}{3}}\]

OpenStudy (acespeedfighter):

@calculusxy so would it be 2 2/3?

OpenStudy (calculusxy):

Yes